# UnsafeBench: Benchmarking Image Safety Classifiers on Real-World and AI-Generated Images

### v2

@@ Introduction
Effectiveness Degradation of Classifiers Trained on Real-World Images Only. Next, we explore in detail why certain classifiers experience degraded performance on AI-generated images, and what characteristics of AI-generated images contribute to it. We select two groups of classifiers that exhibit degraded performances on AI-generated images compared to real-world images. The first group consists of SD_Filter, NFSW_Detector, and NudeNet. These classifiers achieve F1-Scores ranging from 0.650 to 0.833 in categorizing real-world images within the Sexual category. However, the scores drop to 0.596-0.727 when applied to AI-generated images based on Table 4. The second group, including Q16 and GPT-4V, recognizes real-world violent images with F1-Scores of 0.693 to 0.774 but drops to 0.612 to 0.712 for AI-generated violent images. To investigate the reason behind this performance degradation on AI-generated images, we again examine the misclassified examples, including false negatives and false positives from both LAION-5B and Lexica. To characterize images from two sources, we perform KMeans clustering and group these misclassified examples into $K$ clusters, with images from the same cluster sharing similar semantics. Specifically, we utilized the CLIP image encoder to generate image embeddings that capture semantic information. We then apply KMeans clustering with $K=4$ for better visualization purposes. Finally, we create four clusters for each group of images and retrieve the nearest image to each cluster centroid. We display these central images, representing their respective clusters in Figure 5 and Figure 14 in the Appendix, each figure corresponding to one group of classifiers.
@@
Distinct Characteristics in AI-Generated Images. Comparing the misclassified images between real-world and AI-generated ones, we observe two distinct characteristics prevalent in AI-generated images: artistic representation and grid layout. First, images in an artistic representation, [^5] It is an open question whether unsafe content in artistic representation remains unsafe. In this study, we adhere to the criteria from the unsafe image taxonomy and evaluate image safety based on its content and intention, regardless of the representation form. despite showing explicitly unsafe content, can often escape detection of classifiers. For example, among false negatives shown in 5(a), AI-generated images within clusters 2 and 4 show nudity in an artistic style, yet they are misclassified as safe. In contrast, real-world images depicting realistic nudity are rarely misclassified as safe, given that they also widely exist in our annotated dataset. Similarly, among the misclassified images from the Violence category (see 14(a) in the Appendix), AI-generated images within clusters 1, 2, and 3 are also misclassified, which portray artistic representations of violence, e.g., fighting scenes. Second, images in a grid layout from AI-generated ones are frequently misclassified by classifiers. These images are composites made of smaller images, commonly four or nine, arranged in a grid layout. Examples include cluster 1, 3 in 5(a), and cluster 3 in 14(a). Based on these findings, we hypothesize that these distinctive characteristics of AI-generated images, like artistic representation and grid layout, contribute to the reduced effectiveness, particularly for models trained on real-world images, e.g., NSFW_Detector, NudeNet, and Q16.
@@
Case Study on Artistic Representation and Grid Layout. To test our hypothesis, we conduct a case study where we generate image variations with different levels of artistic representation and grid sizes to observe if classifiers can consistently make accurate predictions. To this end, we randomly sample 20 realistic unsafe images from LAION-5B in both Sexual and Violence categories. These images are accurately identified as either sexual content by NSFW_Detector and NudeNet or as violent content by Q16, and these three classifiers have been verified to be trained on real-world images only. For each unsafe image, we then create artistic versions using the shape-preserving image editing function of a text-to-image model, Stable Diffusion [57]. Specifically, we input the image into the model and use the prompt “in the style of oil painting” to generate artistic images. By adjusting the strength parameter, we control the degree of the artistic style applied to the original image. For each image, we produce six artistic versions with the strength parameter ranging from 0.1 to 0.6. Additionally, we generate image variations with different grid sizes by reconstructing an image with $2\times 2$, $3\times 3$, and $4\times 4$ grids, with each smaller image serving as a part of the grid. Figure 6 shows image variations with an example from the Violence category. In total, we collect 40 original images and their 360 variations, including 240 ($40\times 6$) artistic and 120 ($40\times 3$) grid variations. To guarantee that the variations in the artistic representation retain the same content as the original images, we conduct a manual annotation to filter out any variations where the content has been altered. We use the remaining 17 original images and their 153 variations to test NSFW_Detector, NudeNet, and Q16. We present the percentage of correct predictions in Figure 7. The result indicates that with the increasing strength of artistic representation and grid sizes, the classifiers’ effectiveness generally decreases. This confirms the impact of these characteristics found in AI-generated images on the performance of classifiers trained on real-world images only.
@@
SD_Filter 0.999 $\pm$0.001 0.353 $\pm$0.005 $0$$\pm$0.000 0.069 $\pm$0.005 0.355 $\pm$0.001
NSFW_Detector 0.999 $\pm$0.001 0.451 $\pm$0.007 $0$$\pm$0.000 0.325 $\pm$0.009 0.444 $\pm$0.004
@@
Take-Aways. We assess the effectiveness of eight image safety classifiers on real-world and AI-generated images across 11 unsafe categories. Our findings reveal several insights. First, of all the classifiers evaluated, the commercial model GPT-4V stands out as the most effective in identifying a broad spectrum of unsafe content. However, due to its associated time and financial costs, it is impractical to moderate large-scale image datasets. Currently, there is a lack of an open-source image safety classifier that can comprehensively and effectively identify unsafe content. Second, the effectiveness varies significantly across 11 unsafe categories. Images from the Sexual and Shocking categories are detected more effectively, while categories such as Hate require further improvement. For instance, we observe that Neo-Nazi and anti-Semitic symbols can sometimes evade detection by both Q16 and GPT-4V. Finally, we find certain classifiers trained on real-world images experience performance degradation on AI-generated images. Notably, AI-generated images often exhibit unique characteristics, such as artistic representation of unsafe content and unsafe images in a grid layout.
Robustness
Results. We randomly sample 500 images three times that are correctly classified by each classifier and create adversarial examples. Table 5 lists the RA of seven open-source classifiers for four types of perturbations. VLM-based classifiers show the highest robustness. They achieve RAs between 0.563-0.725 on LAION-5B images and 0.634-0.666 on Lexica images, higher than any conventional classifiers. Meanwhile, conventional classifiers are more vulnerable to adversarial attacks, with RAs of 0.299-0.507 on LAION-5B images and 0.290-0.444 on Lexica images. Among them, NudeNet shows the lowest RA (0.290-0.299), revealing that it is the most vulnerable classifier to adversarial attacks.
@@
Classifiers Are More Susceptible to Adversarial Attacks using AI-Generated Images Than Real-World Images. We find that most classifiers, except for InstructBLIP, exhibit lower RAs on AI-generated images compared to their real-world counterparts. Their average RAs show a consistent decrease from the range of 0.299–0.725 to 0.290–0.666, with a maximum drop of 10%. To investigate the reason behind this, we calculate the maximum probability as the confidence score of four conventional classifiers when classifying different groups of images. As shown in Figure 8, successful adversarial examples (denoted in blue) generally have lower confidence scores. This indicates that these examples are closer to the decision boundary and can be more easily perturbed to cross over this boundary compared to non-adversarial examples. When comparing AI-generated and real-world images, AI-generated images tend to have lower confidence scores in the distribution. This potentially contributes to the higher robustness susceptibility for these classifiers.
Interestingly, we also find these classifiers only show evident RA decreases on adversarial perturbations (especially FGSM), and not with Gaussian perturbations. The difference in creating the two types of perturbations is that adversarial perturbations are designed to maximize the training loss, i.e., cross-entropy loss, whereas Gaussian perturbations are not designed to do so. We then analyze classifiers’ cross-entropy loss and their loss change due to the addition of adversarial perturbations, using FGSM as an example. In Figure 15 (in the Appendix), we observe a general trend of higher loss values and loss increase in AI-generated images than real-world ones. This implies that classifiers tend to be more sensitive to the perturbations in AI-generated images, leading to a higher loss increase even with the same amount of perturbation. This also explains why, even if the median confidence scores and loss values are close between AI-generated images and real-world images, such as the NSFW_Detector, the AI image group still makes the classifier more vulnerable to adversarial attacks, because they are more prone to crossing the decision boundary.
PerspectiveVision
@@
The findings from UnsafeBench point out two shortcomings in open-source image safety classifiers, including zero-shot VLMs. First, there is a lack of an open-source image moderation tool that effectively identifies the wide spectrum of unsafe images, particularly in several neglected categories such as hateful content and self-harm. Second, these classifiers show both degraded effectiveness and robustness in detecting AI-generated unsafe images due to distributions that differ from real-world images. With AI-generated content becoming increasingly common in our daily lives, there is an urgent need to develop a comprehensive image moderation tool that accounts for these AI products.
The UnsafeBench dataset serves as a good starting point, as it covers 11 unsafe categories and AI-generated content. Using this dataset, we aim to build a comprehensive moderation tool, PerspectiveVision, that can effectively identify a wide range of unsafe images from the real world and AI. PerspectiveVision is inspired by Google’s Perspective API, a benchmark tool for detecting toxic text in the NLP domain. The API assesses text toxicity across various attributes, such as severe toxicity, hate, and profanity. Similarly, PerspectiveVision is designed to evaluate the safety of a given image and further identify the specific category if the image is deemed unsafe. Users can customize the scope of unsafe images by selecting multiple categories to suit different definitions of unsafe content. The high-level overview is demonstrated in Figure 9.
@@
Table 6: F1-Score of PerspectiveVision models on external evaluation datasets. We use GPT-4V as a baseline for comparison.
In-Distribution Out-of-Distribution (OOD)
Classifier UnsafeBench-Test SMID MultiHeaded-D NudeNet-D Self-Harm Violent Behaviors OOD Average
Linear Probing (CLIP) 0.859 0.365 0.654 0.954 0.801 0.740 0.752
Prompt Learning (CLIP) 0.687 0.623 0.500 0.984 0.971 0.931 0.802
LoRA Fine-Tuning (LLaVA) 0.844 0.557 0.675 0.986 0.948 0.908 0.836
GPT-4V 0.717 0.532 0.577 0.991 0.924 0.921 0.830
@@
The main idea is to build a unified model with VLMs like LLaVA instead of training classifiers separately for different categories. To this end, we fine-tune a VLM, i.e., LLaVA, to identify unsafe images by generating a response that indicates both the safety status and the specific unsafe category. Since fine-tuning the entire LLaVA is computationally intensive, we use Low-Rank Adaptation (LoRA) [40] while training.
Training Dataset Construction. The training dataset is based on our annotated images. The dataset contains triplet elements: image, prompt, and target output. We adopt a similar prompt template as LLaMA-Guard [41] (an LLM-based safeguard), where we define the unsafe image taxonomy with specific unsafe categories and output format. Figure 10 demonstrates an example of the LLaVA fine-tuning dataset. To reduce overfitting, for each annotated image, we randomly remove K irrelevant categories and shuffle the order of the rest categories to build the prompt.
Training Dataset Augmentation. To further improve generalizability, we augment the dataset by label flipping and class balancing. On top of the basic training dataset, for images that are annotated as unsafe, we randomly sample K irrelevant categories to serve as the unsafe image taxonomy in the prompt. We ensure that these images are not unsafe within these categories and then flip the target output to the safe class. This helps the model learn the association between user-defined taxonomy in the prompt and the image label. Finally, we balance the number of samples between two classes by supplementing the smaller class.
Baselines
We employ two common techniques as baselines: linear probing and prompt learning to adapt CLIP to identify unsafe images. Linear probing [54] adds a linear probing layer on top of CLIP, allowing for fine-tuning on specific tasks, and prompt learning [31, 39] refers to a process that adapts CLIP for various downstream tasks by optimizing prompts. For linear probing, we train 11 linear probing layers, each one identifying images from a specific unsafe category. Similarly, we optimize 11 sets of soft prompts using prompt learning.
@@
Evaluation Datasets. We randomly split the UnsafeBench dataset into an 80% training and 20% testing ratio and take the test split as in-distribution evaluation dataset. To assess the model’s generalizability, we further collect multiple external datasets mainly from the training data of conventional classifiers. We regard them as out-of-distribution datasets: SMID [22], MultiHeaded Dataset [12], NudeNet Dataset [16], Self-Harm Dataset [20, 21], and Violent Behavior Dataset [25]. We outline these datasets and detailed experimental setup in Section A.4 and Section A.5 in the Appendix.
Effectiveness Evaluation. Table 6 presents the evaluation results of the PerspectiveVision models across UnsafeBench test set and five OOD datasets. Here, we benchmark GPT-4V against these datasets to establish a high-performing baseline. Among the three methods employed, linear probing (CLIP) and LoRA fine-tuning LLaVA demonstrate a higher fitting level on the UnsafeBench test set, with an overall F1-Score of 0.859 and 0.844. Meanwhile, on external datasets, LoRA fine-tuned LLaVA achieves the highest average F1-Score, 0.836. Although closely followed by GPT-4V, which has an F1-Score of 0.830, our model is notable for its open-source nature, large-scale detection, and fine-grained classification of different unsafe categories. Particularly, it surpasses other models in identifying unsafe images from the MultiHeaded dataset, which contains 800 AI-generated images by different text-to-image models. This improvement can be attributed to the inclusion of a large number of AI-generated images from UnsafeBench in its training data.
@@
Robustness Evaluation. We also test the robust accuracy of LoRA fine-tuned LLaVA against Gaussian and adversarial perturbations. Following the same setup in robustness evaluation, we randomly extract 500 correctly classified images from the UnsafeBench test set and an OOD AI-generated dataset, MultiHeaded-D, and craft adversarial examples with the same perturbation budget. We find that LoRA fine-tuning significantly improves robustness in this image safety classification task. While zero-shot LLaVA shows an average RA of 0.666-0.725 (see Table 5) under four perturbations, fine-tuned LLaVA increases the average RA to 0.952 on UnsafeBench test set and 0.912 on OOD AI-generated images, in Table 7.
Take-Aways. We build a comprehensive image classifier, PerspectiveVision, which identifies unsafe images across 11 categories by fine-tuning LLaVA on our UnsafeBench dataset. The inclusion of diverse AI-generated images not only enhances PerspectiveVision’s effectiveness, particularly with AI-generated images, but also significantly improves its robustness under the same strength of adversarial attacks.
@@
Visual Content Moderation. Moderating visual content is a critical task for both the research community and platform moderators. To mitigate the proliferation of unsafe online images, researchers propose various solutions. For detecting sexual and pornographic images, NudeNet [15] and NSFW_Detector [14] are developed, which are trained on real-world NSFW images. To identify a wider range of unsafe content, Chramowski et al. [60] build Q16 using the prompt learning technique on a dataset containing morally negative/positive images. Additionally, other researchers focus on detecting specific subsets of unsafe images, such as hateful memes [42, 58, 52, 36] and violent protest images [69]. On the commercial front, platform moderators are also actively engaged in proposing moderation solutions. Google’s SafeSearch detection API [3] is capable of identifying unsafe content across five categories: adult, spoof, medical, violence, and racy. Similarly, Microsoft provides an image moderation API [10] that specifically evaluates adult and racy content.

### v3

@@ Introduction
Why are Certain Classifiers Less Effective on AI-Generated Images?
Next, we explore in detail why certain classifiers experience degraded performance on AI-generated images, and what characteristics of AI-generated images contribute to it. We select two groups of classifiers that exhibit degraded performance on AI-generated images compared to real-world images. The first group consists of SD_Filter, NSFW_Detector, and NudeNet. These classifiers achieve F1-Scores ranging from 0.650 to 0.833 in categorizing real-world images within the Sexual category. However, the scores drop to 0.596-0.727 when applied to AI-generated images based on Table 3. The second group, including Q16 and GPT-4V, recognizes real-world violent images with F1-Scores of 0.693 to 0.774 but drops to 0.612 to 0.712 for AI-generated violent images. To investigate the reason behind this performance degradation on AI-generated images, we again examine the misclassified examples, including false negatives and false positives from both LAION-5B and Lexica. To characterize images from two sources, we perform KMeans clustering and group these misclassified examples into $K$ clusters, with images from the same cluster sharing similar semantics. Specifically, we utilize the CLIP image encoder to generate image embeddings and apply KMeans clustering. To determine the optimal K, which maximizes similarity within each cluster while minimizing overlap between different clusters, we apply the Elbow method [3], which calculates the distortion score (indicating how tight the clusters are) and the silhouette score, which measures how well-separated the clusters are. We test K values ranging from 2 to 10, and the optimal K is 4 for the Sexual category (see Appendix 10(a)) and 5 for the Violence category (see Appendix 10(b)). For clearer visualization, we choose K=4. Finally, we create four clusters for each group of images and retrieve the nearest image to each cluster centroid. We display these central images, representing their respective clusters in Figure 4 and Figure 14 in the Appendix, each figure corresponding to one group of classifiers.
@@
Specific Characteristics in AI-Generated Images: Artistic Representation and Grid Layout. Comparing the misclassified images between real-world and AI-generated ones, we observe two specific characteristics prevalent in AI-generated images: artistic representation and grid layout. First, images in an artistic representation, [^4] It is an open question whether unsafe content in artistic representation remains unsafe. In this study, we adhere to the criteria from the unsafe image taxonomy and evaluate image safety based on its content and intention, regardless of the representation form. despite showing explicitly unsafe content, can often escape detection of classifiers. For example, among false negatives shown in 4(a), AI-generated images within clusters 2 and 4 show nudity in an artistic style, yet they are misclassified as safe. In contrast, real-world images depicting realistic nudity are rarely misclassified as safe, given that they also widely exist in our annotated dataset. Similarly, among the misclassified images from the Violence category (see 14(a) in the Appendix), AI-generated images within clusters 1, 2, and 3 are also misclassified, which portray artistic representations of violence, e.g., fighting scenes. Second, images in a grid layout from AI-generated ones are frequently misclassified by classifiers. These images are composites made of smaller images, commonly four or nine, arranged in a grid layout. Examples include cluster 1, 3 in 4(a), and cluster 3 in 14(a). Based on these findings, we hypothesize that these AI characteristics, like artistic representation and grid layout, potentially contribute to the reduced effectiveness, particularly for models trained on real-world images, e.g., NSFW_Detector, NudeNet, and Q16.
A Case Study on Artistic Representation and Grid Layout
@@
To test our hypothesis, we conduct a case study where we generate image variations with different levels of artistic representation and grid sizes to observe if classifiers can consistently make accurate predictions. To this end, we randomly sample 20 realistic unsafe images from LAION-5B in both Sexual and Violence categories. These images are accurately identified as either sexual content by NSFW_Detector and NudeNet or as violent content by Q16, and these three classifiers have been verified to be trained on real-world images only. For each unsafe image, we then create artistic versions using the shape-preserving image editing function of a text-to-image model, Stable Diffusion [57]. Specifically, we input the image into the model and use the prompt “in the style of oil painting” to generate artistic images. By adjusting the strength parameter, we control the degree of the artistic style applied to the original image. For each image, we produce six artistic versions with the strength parameter ranging from 0.1 to 0.6. Additionally, we generate image variations with different grid sizes by reconstructing an image with $2\times 2$, $3\times 3$, and $4\times 4$ grids, with each smaller image serving as a part of the grid. Figure 5 shows image variations with an example from the Violence category. In total, we collect 40 original images and their 360 variations, including 240 ($40\times 6$) artistic and 120 ($40\times 3$) grid variations. To guarantee that the variations in the artistic representation retain the same content as the original images, we conduct a manual annotation to filter out any variations where the content has been altered.
We use the remaining 17 original images and their 153 variations to test NSFW_Detector, NudeNet, and Q16. We present the percentage of correct predictions in Figure 6. The result indicates that with the increasing strength of artistic representation and grid sizes, the classifiers’ effectiveness generally decreases. This confirms the impact of these characteristics commonly found in AI-generated images on the performance of certain classifiers trained on real-world images only, e.g., NSFW_Detector, NudeNet, and Q16.
Takeaways
We assess the effectiveness of eight image safety classifiers on real-world and AI-generated images across 11 unsafe categories. Our findings reveal several insights. First, of all the classifiers evaluated, the commercial model GPT-4V stands out as the most effective in identifying a broad spectrum of unsafe content. Second, the effectiveness varies significantly across 11 unsafe categories. Images from the Sexual and Shocking categories are detected more effectively, while categories such as Hate require further improvement. For instance, we observe that Neo-Nazi and anti-Semitic symbols can sometimes evade detection by both Q16 and GPT-4V. Finally, we find certain classifiers trained on real-world images experience performance degradation on AI-generated images, especially for the Sexual and Violence categories. Notably, AI-generated images exhibit more special characteristics compared to real-world images, such as artistic representation of unsafe content and unsafe images in a grid layout.
Robustness Assessment
Methodology
Adversarial Attacks Against Conventional Classifiers. We evaluate the robustness of classifiers using images perturbed with both random noise (Gaussian noise) and adversarial noise (i.e., adversarial examples) [49, 47]. Adversarial examples ($x_{adv}$) are original inputs ($x$) with optimized perturbations ($\Delta x$), which maximizes the loss ($L(\theta,x+\Delta x,y)$) of a model ($\theta$) and fools the model into making incorrect predictions. The p-norm of the perturbation vector is limited by a small $\epsilon$ to ensure the perturbation is imperceptible.
@@
SD_Filter 0.999 $\pm$0.001 0.353 $\pm$0.005 0.000 $\pm$0.000 0.069 $\pm$0.005 0.355 $\pm$0.001
NSFW_Detector 0.999 $\pm$0.001 0.451 $\pm$0.007 0.000 $\pm$0.000 0.325 $\pm$0.009 0.444 $\pm$0.004
@@
Adversarial Attacks Against VLM-based Classifiers. However, for VLMs, directly solving Equation 1 does not necessarily create adversarial examples that lead to opposite predictions. Unlike binary classifiers, drifting the VLM away from its original predictions, e.g., “the image is safe,” could result in unexpected outputs that are unrelated to the image safety classification task. To address this, we transform the untargeted attack in Equation 1 into targeted attacks (targeting the opposite class), which are equivalent in binary classification settings. For example, if the VLM initially classifies an image as safe, we then optimize the perturbation such that the output moves toward the “unsafe” direction by setting the target output as “unsafe.” Therefore, instead of maximizing the loss between the classifier’s prediction and the original label ($y$), here, we minimize the loss of a VLM between its prediction and the defined target output ($y_{tar}$), as shown in Equation 2. By solving the equation, we update optimized perturbations until the RoBERTa classifier classifies the VLM output as the opposite class from the image label. Using this strategy, we create adversarial examples using FGSM, PGD, and DeepFool on open-source VLM-based classifiers (LLaVA and InstructBLIP). Note that we exclude GPT-4V from this evaluation because crafting adversarial examples requires model gradients, which is not available for GPT-4V.
Consistent Setup. To maintain the same perturbation budget for both conventional classifiers and VLMs, we use the $L_{\infty}$ norm, set $\epsilon$ to 0.01, and also limit the number of optimization iterations to a maximum of 100. We demonstrate the perturbed images and different types of perturbations in Figure 7.
Robustness Result
We randomly sample 500 images three times that are correctly classified by each classifier and create adversarial examples. Table 4 lists the RA of seven open-source classifiers for four types of perturbations. VLM-based classifiers show the highest robustness. They achieve RAs between 0.563-0.725 on LAION-5B images and 0.634-0.666 on Lexica images, higher than any conventional classifiers. Meanwhile, conventional classifiers are more vulnerable to adversarial attacks, with RAs of 0.299-0.507 on LAION-5B images and 0.290-0.444 on Lexica images. Among them, NudeNet shows the lowest RA (0.290-0.299), revealing that it is the most vulnerable classifier to adversarial attacks.
@@
Classifiers Are More Susceptible to Adversarial Attacks using AI-Generated Images Than Real-World Images. We find that most classifiers, except for InstructBLIP, exhibit lower RAs on AI-generated images compared to their real-world counterparts. Their average RAs show a consistent decrease from the range of 0.299–0.725 to 0.290–0.666, with a maximum drop of 10%. To investigate the reason behind this, we calculate the maximum probability as the confidence score of four conventional classifiers when classifying different groups of images. As shown in Figure 15 (Appendix), successful adversarial examples (denoted in blue) generally have lower confidence scores. This indicates that these examples are closer to the decision boundary and can be more easily perturbed to cross over this boundary compared to non-adversarial examples. When comparing AI-generated and real-world images, AI-generated images tend to have lower confidence scores in the distribution. This potentially contributes to the higher robustness susceptibility for these classifiers.
Interestingly, we also find these classifiers only show evident RA decreases on adversarial perturbations (especially FGSM), and not with Gaussian perturbations. The difference in creating the two types of perturbations is that adversarial perturbations are designed to maximize the training loss, i.e., cross-entropy loss, whereas Gaussian perturbations are not designed to do so. We then analyze classifiers’ cross-entropy loss and their loss change due to the addition of adversarial perturbations, using FGSM as an example. In Figure 16 (Appendix), we observe a general trend of higher loss values and loss increase in AI-generated images than real-world ones. This implies that classifiers tend to be more sensitive to the perturbations in AI-generated images, leading to a higher loss increase even with the same amount of perturbation. This also explains why, even if the median confidence scores and loss values are close between AI-generated images and real-world images, such as the NSFW_Detector, the AI image group still makes the classifier more vulnerable to adversarial attacks, because they are more prone to crossing the decision boundary.
Takeaways
In this section, we test the robustness of classifiers against both random and adversarial noises. The robustness evaluation result shows that VLM-based classifiers tend to be more robust compared to conventional classifiers. More importantly, against adversarial examples created with AI-generated unsafe images, classifiers tend to show lower confidence scores and higher loss changes, thus presenting a lower level of robustness. This finding may also connect to adversarial jailbreaking in VLMs, which implies that jailbreaking VLMs with AI-generated images could potentially be more successful. We leave the hypothesis for future work.
Mitigating the Emerging AI Threat
@@
The findings from UnsafeBench point out two shortcomings in open-source image safety classifiers, including zero-shot VLMs. First, tested classifiers present degraded effectiveness on AI-generated unsafe images due to the distribution shift between real-world and AI images, including differences in semantics, noise level, artistic representation, grid layout, etc. Second, these classifiers are more vulnerable to adversarial attacks with AI unsafe images, indicating that AI unsafe images are easier to misclassify with a small amount of noise. Facing these AI threats, we aim to build an image moderation tool with enhanced effectiveness and robustness with AI-generated images.
The UnsafeBench dataset serves as a good starting point, as it covers 11 unsafe categories and AI-generated content. Using this dataset, we aim to build a comprehensive moderation tool, PerspectiveVision, [^5] The name is inspired by Google’s Perspective API, a benchmark tool for detecting toxic text in the NLP domain. which can identify unsafe images based on a user-customized scope of unsafe content. The high-level overview is demonstrated in Figure 8.
@@
Table 5: Effectiveness and generalizability of PerspectiveVision models compared to baselines on both six evaluation datasets. Among these datasets, only the UnsafeBench test (UB-Test) set and the MultiHeaded dataset include AI-generated unsafe images; the others consist of real-world images. The Overall F1 refers to the aggregated F1 score across all six datasets. The Real-World F1 score reflects the performance on real-world unsafe images from six datasets, and the AI-generated F1 score assesses the performance on the AI-generated partition. OOD F1 score measures generalizability on unseen datasets. More specific results for each unsafe category are provided in Appendix Table 12.
Type Model UB-Test SMID NSFW Self-harm Violence MultiHeaded Overall F1 Real-World F1 AI-Generated F1 OOD F1
PerspectiveVision Linear Probing (CLIP) 0.859 0.365 0.949 0.800 0.770 0.654 0.771 0.768 0.782 0.771
Prompt Learning (CLIP) 0.687 0.619 0.982 0.971 0.940 0.508 0.802 0.816 0.604 0.802
LoRA Fine-tuning (LLaVA) 0.844 0.557 0.986 0.948 0.908 0.675 0.836 0.850 0.790 0.840
Classifier Ensemble Q16_NudeNet 0.585 0.642 0.971 0.977 0.929 0.611 0.803 0.840 0.600 0.530
Q16_NSFW_Detector 0.606 0.651 0.986 0.974 0.929 0.601 0.817 0.860 0.600 0.740
Q16_SD_Filter 0.595 0.641 0.928 0.976 0.929 0.619 0.787 0.820 0.610 0.700
Q16_MultiHeaded 0.635 0.652 0.974 0.972 0.938 0.713 0.823 0.850 0.680 0.740
Q16_MultiHeaded_NudeNet 0.625 0.654 0.990 0.978 0.938 0.684 0.822 0.850 0.660 0.750
All Conventional Classifiers 0.607 0.661 0.999 0.982 0.938 0.653 0.817 0.850 0.640 0.750
Commercial Google’s SafeSearch 0.515 0.343 0.997 0.718 0.299 0.522 0.689 0.720 0.540 0.690
Microsoft’s Filter 0.475 0.332 0.988 0.942 0.940 0.539 0.745 0.780 0.480 0.750
GPT-4V 0.717 0.532 0.991 0.924 0.921 0.577 0.830 0.836 0.665 0.830
@@
The main idea is to build a unified model with VLMs instead of training classifiers separately for different categories. To this end, we fine-tune a VLM, i.e., LLaVA, to identify unsafe images by generating a response that indicates both the safety status and the specific unsafe category. Since fine-tuning the entire LLaVA is computationally intensive, we use Low-Rank Adaptation (LoRA) [40] while training.
Training Dataset Construction. The training dataset is based on our annotated images. The dataset contains triplet elements: image, prompt, and target output. We adopt a similar prompt template as LLaMA-Guard [41] (an LLM-based safeguard), where we define the unsafe image taxonomy with specific unsafe categories and output format. Figure 9 demonstrates an example of the LLaVA fine-tuning dataset. To reduce overfitting, for each annotated image, we randomly remove K (1-10) irrelevant categories and shuffle the order of the rest categories to build the prompt.
Training Dataset Augmentation. To further improve generalizability, we augment the dataset by label flipping and class balancing. On top of the basic training dataset, for images that are annotated as unsafe, we randomly sample K irrelevant categories to serve as the unsafe image taxonomy in the prompt. We ensure that these images are not unsafe within these categories and then flip the target output to the safe class. This helps the model learn the association between the user-defined taxonomy in the prompt and the image label. Finally, we balance the number of samples between two classes by supplementing the smaller class.
Baselines. We employ a wide range of techniques and moderation classifiers as baselines. The first group of baselines includes two commonly used techniques: linear probing and prompt learning to adapt CLIP to identify unsafe images. Linear probing [54] adds a linear probing layer on top of CLIP, allowing for fine-tuning on specific tasks, and prompt learning [30, 38] refers to a process that adapts CLIP for various downstream tasks by optimizing prompts. We then train classifiers on the UnsafeBench train set using these two techniques. The second group of baselines includes the tested image safety classifiers as well as their various ensembles, e.g., Q16 + NudeNet, which can cover a wide range of unsafe content. We determine the image is unsafe if any classifier in the ensemble reports it as unsafe. The last group of baselines refers to the commercial moderation models, including GPT-4V, Google’s SafeSearch [4], and Microsoft Content filter [10].
For the experimental setup, we provide the training and evaluation details in Appendix Section A.6.
@@
Evaluation Datasets. We randomly split the UnsafeBench dataset into an 80% training and 20% testing ratio and take the test split as in-distribution evaluation dataset. To assess the model’s generalizability, we further collect multiple external datasets mainly from the training data of conventional classifiers. We regard them as out-of-distribution datasets: SMID [21], MultiHeaded Dataset [11], NudeNet Dataset [15], Self-Harm Dataset [19, 20], and Violent Behavior Dataset [24]. We outline these datasets in Section A.5 in the Appendix.
Effectiveness & Generalizability. Table 5 presents the evaluation results of the PerspectiveVision models across six datasets. Note that several of these evaluation datasets also serve as training data for tested baselines. For example, Q16 is trained on SMID, NudeNet is trained on the NSFW dataset, and the MultiHeaded classifier is trained on the MultiHeaded dataset. Therefore, directly comparing model performance on each individual dataset may not be fair. To address this, we focus on the aggregated performance across all datasets. Among all evaluated models, the fine-tuned LLaVA achieves the highest overall F1 score of 0.836, across six evaluation datasets. Notably, it also obtains the highest F1 score on AI-generated images, reaching 0.796. This improvement can be attributed to the inclusion of a large number of AI-generated images from UnsafeBench in its training data.
To evaluate generalizability, we calculate the out-of-distribution (OOD) F1 score, i.e., performance on unseen datasets. For example, for PerspectiveVision, all datasets excluding the UnsafeBench test set are treated as unseen datasets; for Q16 and NudeNet, datasets other than SMID and NSFW are considered OOD. As shown in Table 5, the fine-tuned LLaVA achieves the highest OOD F1 score, indicating strong generalization capability.
To conclude, the fine-tuned LLaVA presents the highest effectiveness and generalizability on six evaluation datasets, serving as the top-performing checkpoint in PerspectiveVision models.
@@
Robustness. We also test the robust accuracy of LoRA fine-tuned LLaVA against Gaussian and adversarial perturbations. Following the same setup in robustness evaluation, we randomly extract 500 correctly classified images from the UnsafeBench test set and an OOD AI-generated dataset, MultiHeaded-D, and craft adversarial examples with the same perturbation budget. We find that LoRA fine-tuning significantly improves robustness in this image safety classification task. While zero-shot LLaVA shows an average RA of 0.666-0.725 (see Table 4) under four perturbations, fine-tuned LLaVA increases the average RA to 0.952 on the UnsafeBench test set and 0.912 on OOD AI-generated images, in Table 6.
Takeaways. We build a comprehensive image classifier, PerspectiveVision, which identifies unsafe images across 11 categories by fine-tuning LLaVA on our UnsafeBench dataset. The inclusion of diverse AI-generated images not only enhances PerspectiveVision’s effectiveness, particularly with AI-generated images, but also significantly improves its robustness under the same strength of adversarial attacks.
@@
Visual Content Moderation. Moderating visual content is a critical task for both the research community and platform moderators. To mitigate the proliferation of unsafe online images, researchers propose various solutions. For detecting sexual and pornographic images, NudeNet [14] and NSFW_Detector [13] are developed, which are trained on real-world NSFW images. To identify a wider range of unsafe content, Schramowski et al. [60] build Q16 using the prompt learning technique on a dataset containing morally negative/positive images. Additionally, other researchers focus on detecting specific subsets of unsafe images, such as hateful memes [42, 58, 52, 35] and violent protest images [66]. On the commercial front, platform moderators are also actively engaged in proposing moderation solutions. Google’s SafeSearch detection API [4] is capable of identifying unsafe content across five categories: adult, spoof, medical, violence, and racy. Similarly, Microsoft provides an image moderation API [10] that specifically evaluates adult and racy content.